# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v4

@@ V-A2 Bridging the gap from simulation to reality
Creating a reliable simulation environment and conducting RL training is a complex task. Advanced simulators are essential for training policies using full-order dynamics, which better represent the complex interactions robots face in the real world. Simulators like Isaac Gym [81], RoboCup3D [82], OpenAI Gym [83], and MuJoCo [84] are widely used to train policies in environments that closely mimic physical conditions. These tools help minimise the sim-to-real gap, improving the success of policy transfer to real robots. Further details on these simulators can be found in Appendix B. However, the process of sim-to-real encounters significant challenges due to the approximations and simplifications made in simulations. The process of sim-to-real faces significant challenges due to the approximations and simplifications made in simulations. Two primary issues arise: (1) Inaccurate dynamic models, contribute to discrepancies between simulation environments and real-world dynamics. They often simplify or generalise complex physical interactions, leading to a gap in performance when transitioning to reality. (2) The real environment is more diverse and complex, with varying terrain, sensor noise, and unpredictable conditions not captured in simulations. Numerous sim-to-real frameworks [43, 85, 86] have shown high efficiency and performance. Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments.
To bridge the sim-to-real gap, several approaches have been developed, including dynamics randomisation [44], system identification [87, 88], periodic reward composition [49], learned actuator dynamics [87, 89], regulation feedback controllers [90], and adversarial motion priors [27, 26]. In domain randomisation, there are two main methods for training policies: (1) end-to-end training using robot measurement histories, and (2) policy distillation, where an expert policy (e.g., teacher-student [49]) guides a student policy that learns from internal sensory feedback, as seen in techniques like RMA [85].
Future research in bridging the sim-to-real gap will focus on enhancing the transferability and robustness of policies across diverse environments, ultimately aiming for zero-shot transfer, where no real-world fine-tuning is required. This will not only involve improving hardware but also making simulation environments more reflective of real-world conditions. Additionally, advancing sophisticated vision-based perception and multimodal sensory integration will be crucial in further minimising the sim-to-real gap, making policies more adaptable to complex and unpredictable environments.
V-A3 Integrating foundation models
Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots, as discussed in Section IV. Their integration into robotics extends beyond decision-making, offering new opportunities to enhance control, adaptability, and learning efficiency. One particularly promising direction is their use in RL, not merely for planning but for shaping the learning process itself—most notably in the design of reward functions, which remains a core challenge in bipedal locomotion tasks.
The design of effective reward functions is a fundamental challenge in DRL for bipedal robots. While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem. These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming. Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
LLMs can enhance RL by guiding policy optimisation and refining reward functions. A key challenge in DRL is reward function design, which typically requires manual tuning. LLMs offer a solution by dynamically adjusting reward signals based on linguistic descriptions of task success.
For example, Kumar et al. [91] demonstrate how LLMs refine reward functions by incorporating linguistic feedback, improving hierarchical RL frameworks. The language model processes HL instructions, which are then translated into lower-level RL tasks, simplifying complex skill acquisition and human-robot interactions. Conversely, end-to-end approaches such as AnyBipe [92] use LLMs to autonomously adjust reward functions and control strategies, reducing manual intervention. This enables more autonomous learning, allowing the model to self-optimise in diverse environments and handle complex terrain navigation.
Other implementations of foundation models in RL have been comprehensively reviewed in [60]. Future research should further explore the synergy between LLMs and DRL, with a particular focus on adaptive reward shaping and self-optimising control policies. Advancing these methodologies could pave the way for more autonomous, self-learning humanoid robots, capable of dynamically refining their decision-making processes and adapting seamlessly to diverse environments with minimal human intervention.
V-A4 Vision-based learning framework
The vision-based learning framework plays a crucial role in advancing the unified framework for bipedal robots. Integrating vision into the unified system enhances the robot’s ability to perceive and interpret its environment, making it capable of handling more complex and dynamic tasks, such as navigating rough terrains or interacting with objects in real-time. Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [46, 53]. This can be attributed to several factors: (1) Simplified Problem: Blind locomotion reduces task complexity, focusing directly on motor control strategies rather than processing visual inputs. (2) Computation and Training Complexity: Integrating vision requires significant computational resources and robust algorithms for real-time processing, terrain identification, and object detection. (3) Advanced Control Without Vision: Previous studies have already demonstrated robust blind locomotion using proprioceptive data, such as joint angles, IMU, with vision typically enhancing precision in locomotion [52] or path planning [57].
However, recent advancements in computing power and computer vision technology—such as more efficient real-time visual processing algorithms and increasingly powerful hardware—have significantly accelerated the integration of vision-based learning into robotics. This shift has led to a surge in the popularity of vision-based approaches, allowing robots to more effectively interpret and adapt to complex environments, making them a central component in modern frameworks.
Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal robots [29, 52, 36]. Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
  • Height scanner mapping: This approach, evaluated in works like [52], involves using height maps generated by scanners to inform locomotion strategies. These maps provide detailed topographical data, allowing robots to plan steps on uneven or obstructed surfaces more effectively.
  • Direct vision inputs: Directly utilising inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [57, 93]. Although previous studies like [57] have integrated visual navigation by feeding visual information to a HL planner, the potential of direct visual inputs to RL policies has not been fully explored.
Enhancing the capability of bipedal robots to directly interpret and utilise visual data without intermediary processing can revolutionise their adaptability and efficiency in real-world scenarios. The exploration of direct vision inputs to RL-based policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
V-B Insights from quadruped robots
@@
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [94, 95], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [96]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [97], and sophisticated acrobatic motions are achieved through imitation learning [98].
@@
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [99] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [59, 100] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs a Model Predictive Control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [76] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
@@
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway. These include mimicking real animals through motion capture data and imitation learning [101, 102], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [103, 104] or by using the robots’ legs [105]. Notably, the research presented in [104] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.

### v5

@@ IV-B Reality gap: From simulation to the physical world
Simulators like Isaac Gym [67], RoboCup3D [76], OpenAI Gym [77], and MuJoCo [78] detailed in Appendix B are widely used to train policies that closely mimic real-world physical conditions. These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [46, 79, 80] have demonstrated efficient and high-performance results. Despite these advancements, a significant gap persists between simulation and reality, exacerbated by the approximations made in simulation and the unpredictability of physical environments.
IV-C Safety-critical Locomotion
Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [1, 15, 16]. This includes ensuring the robot’s own integrity to prevent costly damage, as well as guaranteeing the safety of the surrounding environment and any humans within it. While many existing frameworks have demonstrated impressive locomotion skills, they often prioritize performance over these safety considerations. This creates a critical barrier that separates success in controlled laboratory settings from reliable operation in the unpredictable real world.
  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) like joint angles and IMU data [47, 49, 32], creating a major safety risk. Lacking external perception, these robots cannot anticipate obstacles, slopes, or slippery surfaces, making them purely reactive and highly prone to failure. Despite these significant safety drawbacks, this approach is often adopted for several key reasons. Omitting vision simplifies the control problem to pure motor skills and avoids the computational complexity of real-time visual processing. Furthermore, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [55] or path planning [60], rather than a core requirement for basic stability.
  • Lack of physical constraint satisfaction: Many DRL frameworks lack built-in mechanisms to guarantee physical constraint satisfaction. This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules, for example Safe-RL on humanoids [81], hybrid DRL with identified low-dimensional safety models [82], footstep-constrained DRL policies [54], and reactive DRL steppers operating under feasibility constraints on uneven terrain [72]. This limitation makes it difficult to prevent the robot from exceeding joint limits, applying excessive torques, or causing self-collisions, particularly when reacting to unexpected events. This is a key area where constrained RL could be applied.
In summary, the pursuit of performance in DRL has often sidelined critical safety issues. The prevalence of blind policies that cannot anticipate environmental hazards, combined with the lack of inherent mechanisms to enforce physical constraints, creates a significant risk that hinders real-world deployment. While these challenges are significant, they also define a clear path forward. The following section on Future Directions and Opportunities will explore specific research avenues, such as vision-based learning and safe reinforcement learning, aimed directly at overcoming these safety barriers and enabling the development of truly robust and reliable bipedal robots.
Figure 4: Towards a Unified Framework: This figure illustrates the logical progression from current DRL frameworks to future unified systems. It identifies the current limitations of existing end-to-end and hierarchical approaches, which motivate the exploration of specific Future Pathways. These pathways inform the design of two proposed conceptual models (i) Multi-Layered Adaptive Model (MLAM) and (ii) Bipedal Foundation Model (BFM) which represent potential blueprints for achieving a generalist, unified framework.
V Future Directions and Opportunities
Following the analysis of the surveyed frameworks and their limitations, this section outlines a path forward for DRL-based bipedal locomotion by exploring both direct research avenues and emerging opportunities. We begin in Section V-A by detailing research directions that directly respond to the challenges identified in Section IV. Building on this foundation, Section V-B broadens the scope to explore synergistic opportunities from related fields, such as loco-manipulation and the application of foundation models. These discussions culminate in Section V-C, where we propose two conceptual models for a unified framework that represent the future evolution of the end-to-end and hierarchical paradigms.
V-A Pathways for bipedal locomotion
V-A1 Multi-skill learning
A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [63, 83, 35]. This research direction focuses on enabling robots to acquire, adapt, and deploy a broad and varied repertoire of motor skills, allowing them to handle unforeseen situations and operate effectively in unstructured environments.
To achieve such versatility, researchers are pursuing several HL pathways, which can be broadly categorized into structured and holistic approaches. The structured approach focuses on explicit decomposition. A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, a HL planner generates reference trajectories, while a LL DRL controller executes them robustly [59, 56], has shown in Section III. Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [26]. A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single compact generalist policy [84].
V-A2 Traversing challenging environments
The goal of versatile skill learning is to enable bipedal robots to traverse challenging, human-centric environments where their unique form has an advantage. Validating capabilities on such terrains serves a crucial dual purpose. It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [49, 52, 63]. More critically, it benchmarks precision on treacherous paths such as stepping stones that demand exact foot placement [33, 19, 72]. These environments are the ultimate test of both a robot’s skill repertoire and its control fidelity.
V-A3 Efficient learning and reward shaping
As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [40, 32]. Addressing this bottleneck is a crucial research frontier that calls for both more efficient algorithms and more robust reward designs.
To mitigate sample inefficiency for complex skills, several research pathways are being actively explored. A primary strategy is to leverage prior data rather than learning entirely from scratch. Leveraging prior knowledge provides strong guidance and reduce unsafe exploration by anchoring policies to feasible motion patterns [26, 45, 59]. Curriculum learning further organizes training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [53, 51, 39].
Complementing advances in algorithms is the design of effective and robust rewards. Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [68, 26]. Phase aware objectives are well established for cyclic gaits such as walking [32], whereas reward design for nonperiodic skills such as jumping is less standardized and often task specific [48]. Promising directions reduce manual effort by adding higher level guidance, including event based terms, goal conditioned objectives, and kinematic reference tracking [54, 28, 50]. Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand crafted objectives with implicit ones inferred from demonstrations [85]. Together, these directions seek to minimize skill specific tuning and improve the transferability and reliability of learned locomotion policies.
V-A4 Motion retargeting
As human-like agents, bipedal robots, especially humanoids, have the unique advantage of a morphology that is similar to our own. This presents a significant opportunity: the potential to learn from vast libraries of human motion data. While large-scale datasets like AMASS [86] and Motion-X [87] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [29]. Therefore, Motion Retargeting emerges as a critical component to bridge this gap. The challenge of this pathway is not merely to transfer human movements to the robot, but to generate trajectories that are both high-quality in their stylistic fidelity and rigorously physically feasible, adhering to the robot’s unique dynamics and constraints. Successfully developing these retargeting methods provides a scalable solution for accessing the data needed to train the natural and versatile generalist policies of the future.
V-A5 Bridging the gap from simulation to reality
Strategies to bridge the sim-to-real gap generally follow two main philosophies. The first aims to train policies that are robust enough to tolerate the inevitable mismatch between simulation and reality. The second focuses on minimizing the gap itself by making the simulator a more faithful replica of the physical world.
The first approach seeks to reduce the discrepancy by improving the simulation’s fidelity. This is often achieved through system identification (SI), where real-world robot data is used to fine-tune the simulator’s parameters to create a more accurate ”digital twin” [74, 88]. This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [74, 89]. Other methods, such as designing specialized feedback controllers [90], also contribute by making the system less sensitive to the remaining modeling errors.
In contrast, the second philosophy accepts that simulations will always be imperfect and instead focuses on creating highly adaptive, robust policies. The primary method here is DR, which forces a policy to generalize by training it across a wide range of simulated physical variations. Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [79], or via policy distillation, where a privileged “teacher” guides a “student” policy [52] to have a knowledge of unknown information like friction . Additionally, techniques like adversarial motion priors [31, 30] are used to ensure the learned behaviours are not just robust but also physically plausible.
Looking ahead, the ultimate goal remains achieving reliable zero-shot transfer, where no real-world fine-tuning is needed. Progress will depend on the co-development of higher-fidelity simulations, improved hardware, and more robust control policies inherently capable of handling real-world unpredictability. The synergy of these advancements will be crucial in finally closing the sim-to-real gap.
V-A6 Perception-conditioned locomotion
ntegrating exteroceptive sensors like cameras and LiDAR enables bipedal robots to proactively plan footsteps, avoid obstacles, and adapt to upcoming terrain. This shift from reactive to anticipatory control is essential for navigating unstructured real-world environments.
The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [33, 55, 39]. In contrast, LiDAR is an active sensing method that generates precise 3D point clouds of the terrain. While vision provides richer data but is sensitive to lighting, LiDAR offers robust geometric measurements without visual detail.
Based on this sensory data, current research is exploring two primary pathways for processing perceptual information for control. The first involves creating an intermediate geometric representation, such as a height map from scanners [55]. This provides the policy with structured topographical data for effective footstep planning. The second is a more end-to-end approach, which utilizes direct vision inputs like RGB or depth images as inputs to the RL policy for real-time decision-making [60, 91]. The former offers interpretability, while the latter promises more nuanced, reactive behaviours learned directly from raw perception.
Future progress requires advancing both pathways: building richer, semantic world representations and improving the efficiency of direct perception-to-action policies. Solving the underlying challenges of real-time processing and the perceptual sim-to-real gap will be crucial for enabling truly adaptive locomotion in complex, real-world scenarios.
V-A7 Constrained learning
While the previously discussed pathways focus on enhancing a robot’s capabilities, a parallel and equally critical frontier is ensuring that these capabilities are exercised safely and reliably. To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [81, 82].
Soft constraints encourage desirable behaviour and penalize undesirable behaviour without forbidding it. They are well suited for preferences or efficiency goals, for example minimizing energy use, limiting peak torques, or promoting smooth motion [81]. Hard constraints are inviolable rules that prevent catastrophic failures. They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [54, 90]. A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [92].
In practice, a robust and trustworthy bipedal robot will likely combine both ideas. Soft constraints help a policy learn efficient and natural gaits, while hard constraints guarantee that it will not take catastrophic actions. This combination supports the transition from systems that are merely capable in laboratory settings to agents that are reliable, predictable, and safe for real-world deployment.
V-B Opportunities
V-B1 Leveraging foundation models for locomotion learning
The recent rise of Foundation Models (FMs), such as Large Language Models (LLMs) and Vision-Language Models (VLMs), presents a transformative opportunity for bipedal locomotion. Their powerful reasoning capabilities are unlocking new approaches that go beyond traditional control methods, primarily by enabling sophisticated HL task planning and by providing novel solutions to shape the learning process itself, particularly in automated reward design.
As HL planners, FMs provide a reasoning engine that can bridge the gap between abstract human goals and LL motor execution. They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for a LL policy to follow. This has been demonstrated effectively in legged robotics where VLMs process raw sensory data to pass structured commands to motor controllers [93], creating a seamless link between strategic planning and physical action.
Furthermore, FMs create a significant opportunity to overcome one of the most persistent bottlenecks in DRL: reward design. Instead of tedious manual tuning, LLMs can dynamically generate or refine reward functions based on linguistic descriptions of task success. Research has shown that LLMs can translate human feedback into reward adjustments [94] or even autonomously adjust rewards and control strategies to self-optimize for diverse terrains [95], drastically reducing human intervention.
The foremost opportunity lies in the deeper synergy between these roles. The integration of the HL symbolic reasoning of FMs with the LL, real-time control of DRL could create a new class of highly adaptive and flexible robots. As this rapidly evolving field progresses, as reviewed in [96], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
V-B2 Loco-manipulation tasks
While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it can’t physically interact with its environment. The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction. Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [42].
However, realizing this opportunity is a significant challenge. Early studies, such as a ‘box transportation’ framework [97], often rely on inefficient, multi-policy solutions that lack visual perception. Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [98, 99].
These difficulties create significant research opportunities. One such opportunity lies in exploring hierarchical control approaches. By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [56].
Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution. Using techniques like curriculum learning and imitation from human motion capture data [51, 39, 35, 100, 31, 29], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
V-B3 Insights from quadruped robots
@@
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [101, 102], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [103]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [104, 105], and sophisticated acrobatic motions are achieved through IL [106].
@@
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [107] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [62, 108] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs a MPC to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots module [109] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
@@
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway. These include mimicking real animals through motion capture data and IL [110, 111], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [112, 113] or by using the robots’ legs [114]. Notably, the research presented in [113] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
@@
V-C Conceptual Models for Unified Frameworks
Motivated by our survey and the current state of the art, we propose two conceptual models, intended as reference designs, toward a unified locomotion framework. They build on end-to-end and hierarchical paradigms and offer complementary routes to scalable, generalizable architectures, rather than fully realized systems.
  • Bipedal Foundation Models (BFMs): large-scale, pre-trained models that map perception directly to action through representation learning. Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
  • Multi-Layer Adaptive Frameworks (MLAMs): modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs. MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
In the following sections, we will analyze each of these conceptual models in detail, evaluating their respective strengths and challenges in the pursuit of a unified framework.
V-C1 Bipedal Foundation Models
Inspired by Robot Foundation Models (RFMs) [96, 115], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion. A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments. Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots. Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [116].
The proposed BFM paradigm would operate in two stages. First, IL on diverse datasets would establish the generalisable foundation. Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot. The potential of this approach is highlighted by recent works, with frameworks like FLaRe [117] enhancing generalisation for long-horizon tasks, MOTO [118] enabling effective offline-to-online adaptation from images, and AdA [119] demonstrating in-context adaptation to novel environments. Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
However, realizing the BFM concept for bipeds presents significant challenges. The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [117, 119]. Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
V-C2 Multi-layered Adaptive Models
As a complementary path to BFMs, we propose the concept of MLAMs. Rather than relying on large-scale pre-training, this conceptual framework would adopt a modular, hierarchical approach. The idea is to extend conventional hierarchical frameworks (discussed in Section III) with explicitly adaptive layers, allowing for the dynamic composition of specialized policies. The core principle of this concept would be modularity, enabling each layer to be independently optimized or replaced and providing interpretable outputs at each stage.
A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer leverages large pre-trained model such as LLMs and VLMs [120, 121] to parse commands into sub-tasks. For instance, Vision-Language Model Predictive Control [93] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
The mid-level planner selects or synthesises motions via learned motion libraries [58, 44] or DRL-based planners [122]. The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [123], adaptive tracking controllers for whole-body tracking [59], and imitation-based skills such as kicking and dancing [26], by utilizing RL, IL, or model-based methods. This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [123].
However, realizing the MLAM concept would introduce challenges distinct from BFMs. Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [109]. Additionally, the computational latency incurred by HL reasoning modules like LLMs [123] would pose limitations for tasks needing rapid reactions.